# The orbital evolution of the tidally stripped star and disk-driven stable mass transfer for QPEs in GSN 069

### v3

@@
11institutetext: Department of Physics, Huazhong University of Science and Technology, Wuhan 430074, China
The origin of the quasi-periodic eruptions (QPEs) is possibly mass loss at the periastron of a body moving around the supermassive black hole (SMBH) in a high eccentric orbit. Such a tidally stripped star is expected to radiate gravitational wave thereby leading to shrinkage of the periastron distance, and thus will eventually be disrupted by the SMBH in the previous studies. This scenario predicts a gradually increasing mass transfer, contradicting the observed decay in the intensity of the QPEs in GSN 069. In this paper, we first revisited the orbital evolution of the stripped star. The effect of the mass transfer finally dominates the orbital evolution, resulting in the stripped star finally escaping the SMBH rather than being disrupted by it. Then we suggested a model of a tidal stripped WD moving inside an accretion disk for QPEs in GSN 069. The drag force by the disk can effectively reduce mass transfer and thus explain the observed decay in the intensity of the QPEs in GSN 069. The disk is likely a fallback disk of the tidal disruption event in GSN 069. Considering the evolution of its accretion rate, the increase in the intensity of the latest eruption can also be explained.
The orbital evolution of the tidally stripped star and disk-driven stable mass transfer for QPEs in GSN 069
Di Wang
@@ 1 Introduction
Due to gravitational wave radiation decreasing the periastron distance, the mass loss of the tidally stripped WD gradually increases until finally the WD is tidally disrupted by the SMBH (Zalamea et al., 2010; Wang et al., 2022). However, the intensity of QPEs in GSN 069 persistently decays (Miniutti et al., 2023b), which is inconsistent with the previous prediction. King (2022) points out that the mass transfer stability requires the angular momentum transfer from the disk to the orbiting star, he suggested that orbital resonance with the disk possibly could accomplish this, but further discussion was lacking.
GSN 069 also has a tidal disruption event(TDE) that has lasted more than a decade (Shu et al., 2018; Sheng et al., 2021). This TDE was recently found to have re-brightened, and while it was brightening, the QPEs disappeared, probably due to being covered by radiation from the TDE Miniutti et al. (2023b). When radiation from the TDE faded again, the QPEs reappeared and were stronger than in the last outburst before they disappeared (Miniutti et al., 2023a). Such behavior implies that the evolution of mass transfer may be related to the accretion rate of the accretion disk.
@@
While the origin of QPEs is still debated, there are various other models (Chen et al., 2022; Zhao et al., 2022; Wang et al., 2022; Krolik & Linial, 2022; Lu & Quataert, 2023; Xian et al., 2021; Suková et al., 2021; Linial & Sari, 2023; Linial & Metzger, 2023). In this paper, we focus on the stripped WD model in a high eccentric orbit for GSN 069 (King, 2020, 2022). We calculated the secular effect of drag force by the disk on the orbit for GSN 069. We found that the drag force can transfer the angular momentum of the disc to the stripped WD, and then stabilize the mass transfer. This effect is strongly dependent on the accretion rate, which can be estimated from the luminosity of the TDE. The derived change in mass transfer rate is in very good agreement with the observed intensity decay of QPEs in GSN 069. The increase in the intensity of the QPEs after the TDE re-brightened can also be explained by this model.
@@ 2.5 Comparison with the observation of QPEs
QPEs from GSN 069 were rapidly decaying and disappeared as TDE brightened again (Miniutti et al., 2023b). Then, Miniutti et al. (2023a) found that it had reappeared and was brighter than before. Orbital evolution under mass transfer and gravitational radiation does not explain these variations, the $\Delta M$ always increased rapidly or is constant, as shown in Fig. 1.
@@ 3.2.1 Stable mass transfer
Here we adopt the results in Miniutti et al. (2023b). Assuming radiation efficiency $\eta=0.1$, the accreted mass is about $3.7\times 10^{-8}\mathrm{\leavevmode\nobreak\ M_{\odot}}$, corresponding to $\dot{M}/M\approx-1.7\times 10^{-4}\mathrm{yr}^{-1}$ and $\beta\approx 0.5005$ with $M=0.21\mathrm{\leavevmode\nobreak\ M_{\odot}}$. The QPEs intensity is roughly halved per year, corresponding to $\dot{\Delta M}/\Delta M=-0.5\mathrm{\leavevmode\nobreak\ yr}^{-1}$ in Eq. (2), which requires $\dot{\beta}/\beta=-2\times 10^{-4}\mathrm{\leavevmode\nobreak\ yr}^{-1}$.
The change of $\beta$ induced by mass transfer is $\langle\dot{\beta}/\beta\rangle_{\dot{M}}\approx 1.1\times 10^{-4}\mathrm{\leavevmode\nobreak\ yr}^{-1}$ with $\zeta=-1/3$. The bolometric luminosity of TDE is about $10^{44}\mathrm{\leavevmode\nobreak\ erg/s}$ corresponding to $\sim 1.6\dot{M}_{Edd}$. Assuming $e=0.95$, it predicts $\langle\dot{\beta}/\beta\rangle_{disk}\approx-3.5\times 10^{-4}\mathrm{\leavevmode\nobreak\ yr}^{-1}$ as shown in Fig. 6. Both of them give $\dot{\beta}/\beta\approx-2.4\times 10^{-4}\mathrm{\leavevmode\nobreak\ yr}^{-1}$, which is consistent with the required value. Consequently, the hydrodynamic drag force can explain the decaying of the intensity of QPEs from GSN 069.
@@ 3.2.2 History and future of evolution
As the accretion rate decreases, drag force by the disk begins to dominate $\dot{\beta}$, and at that point the intensity of the QPEs begins to decrease. As the TDE rebrighten, QPEs disappeared, and the increased accretion rate led to a weaker drag force and thus an increase in the intensity of the QPEs. It is consistent with the recent observation, which shows that QPEs in GSN 069 reappeared with a higher intensity than before they disappeared (Miniutti et al., 2023a).
As the accretion rate continues to decrease, the effect of the drag force enhances and the intensity of QPEs will keep decreasing. After switching from radiation pressure dominance to gas pressure dominance in the inner region of the disk, the drag force decreases with decreasing accretion rate. When $\dot{\beta}$ is dominated by gravitational wave radiation, the mass transfer begins to enhance. When $\dot{\beta}$ is dominated by mass transfer, the unstable mass transfer leads to WD escape from the SMBH.
@@ 4.3 On the other QPEs
The TDE in GSN 069 makes it unique. On the one hand, it may be the only source whose accretion disk is coplanar with the orbit of the stripped star. On the other hand, the accretion rate of its disk is dominated by the TDE and therefore evolves rapidly. These two points lead to a complicated orbital evolution, making the QPEs in GSN 069 an excellent sample to use to study the interaction between the disk and the star moving inside it.
@@ 5 Summary
The TDE in GSN 069 may be a disruption of the common envelope before the WD was captured, so the orbiting WD is likely moving inside the fallback disk. The hydrodynamic drag force by the disk can effectively reduce $\beta$ and stabilize the mass transfer, which can explain the observed decay in the intensity of QPEs. After the TDE re-brightened, the mass transfer increases due to the weakening of the drag force caused by the reduced density of the inner zone of the accretion disk, which is consistent with the observed increase in the intensity of the latest QPEs.
As the accretion rate of the fallback disk gradually decreases, the drag force by the disk will decrease until it is negligible. Then the mass transfer will start to be unstable eventually the WD will escape the SMBH. Continued observations of GSN 096 will help us to understand the interaction of the disk with the star inside it and to test our model.
@@ Acknowledgements.
This work was supported by the National Key Research and Development Program of China (No. 2020YFC2201400) and National SKA Program of China (2020SKA0120300).

### v4

@@
Abstract
Context. The origin of the quasi-periodic eruptions (QPEs) is possibly mass loss at the periastron of a body moving around the supermassive black hole (SMBH) in a high eccentric orbit. Such a tidally stripped star is expected to radiate gravitational wave thereby leading to shrinkage of the periastron distance, and thus will eventually be disrupted by the SMBH in the previous studies.
Aims. This scenario predicts a gradually increasing mass transfer, contradicting the long term evolution of the observed intensity of the QPEs in GSN 069.
Methods. In this paper, we first revisited the orbital evolution of the stripped star. Then we suggested a model of a tidal stripped WD moving inside an accretion disk for QPEs in GSN 069.
Results. We found the effect of the mass transfer finally dominates the orbital evolution, resulting in the stripped star finally escaping the SMBH rather than being disrupted by it. The drag force by the disk can effectively reduce mass transfer and thus explain the observed long term evolution in the intensity of the QPEs in GSN 069. The disk is likely a fallback disk of the tidal disruption event in GSN 069. Considering the evolution of its accretion rate, the increase in the intensity of the latest eruption can also be explained.
@@ 1 Introduction
Due to gravitational wave radiation decreasing the periastron distance, the mass loss of the tidally stripped WD gradually increases until finally the WD is tidally disrupted by the SMBH (Zalamea et al., 2010; Wang et al., 2022). However, the intensity of QPEs in GSN 069 shows no sign of increasing (Miniutti et al., 2023b, a), which is inconsistent with the previous prediction. King (2022) points out that the mass transfer stability requires the angular momentum transfer from the disk to the orbiting star, he suggested that orbital resonance with the disk possibly could accomplish this, but further discussion was lacking.
GSN 069 also has a tidal disruption event(TDE) that has lasted more than a decade (Shu et al., 2018; Sheng et al., 2021). This TDE was recently found to have re-brightened, and while it was brightening, the QPEs disappeared (Miniutti et al., 2023b). When radiation from the TDE faded again, the QPEs reappeared and were stronger than in the last outburst before they disappeared (Miniutti et al., 2023a). Such behavior implies that the evolution of mass transfer may be related to the accretion rate of the accretion disk.
@@
While the origin of QPEs is still debated, there are various other models (Chen et al., 2022; Zhao et al., 2022; Wang et al., 2022; Krolik & Linial, 2022; Lu & Quataert, 2023; Xian et al., 2021; Suková et al., 2021; Linial & Sari, 2023; Linial & Metzger, 2023; Franchini et al., 2023). In this paper, we focus on the stripped WD model in a high eccentric orbit for GSN 069 (King, 2020, 2022). We calculated the secular effect of drag force by the disk on the orbit for GSN 069. We found that the drag force can transfer the angular momentum of the disc to the stripped WD, and then stabilize the mass transfer. This effect is strongly dependent on the accretion rate, which can be estimated from the luminosity of the TDE. The derived change in mass transfer rate by the disk drag has the ability to suppress other effects that increase mass transfer, thus explaining the long-term evolution of the observed intensity of QPEs. The increase in the intensity of the QPEs after the TDE re-brightened can also be explained by this model.
@@ 2.1 Mass transfer
Here $\langle\dot{J}/J\rangle\approx 0$, because only very little angular momentum is removed by central SMBH via accretion(King, 2022).
@@ 2.5 Comparison with the observation of QPEs
Before QPEs from GSN 069 disappeared, its peak intensity in 0.4–1 keV decays dramatically (Miniutti et al., 2023b). However, the bolometric luminosity by assuming absorbed black-body spectrum was almost constant for each observation, except for the last one, where it was reduced by roughly half (Miniutti et al., 2023a). If the bolometric luminosity is positively correlated with the mass loss of the star at the periastron, such result implies that $\beta$ is constant and then gets smaller.
Assuming the QPEs is caused by the mass transfer via Roche-lobe overflow, the orbiting star should be a WD for GSN 069 (King, 2020). As shown in Fig. 1, the mass loss for WD grows rapidly except in the final stage of evolution, when the orbital period also grows rapidly. There was no significant change in the period of QPEs in GSN 069. So a mechanism for stabilizing mass transfer is required to explain the long-term evolution of QPEs, and the drag of a gaseous disk may achieve this.
@@ 3.2.1 Stable mass transfer
Here we adopt the results in Miniutti et al. (2023b). Assuming radiation efficiency $\eta=0.1$, the accreted mass is about $3.7\times 10^{-8}\mathrm{\leavevmode\nobreak\ M_{\odot}}$, corresponding to $\dot{M}/M\approx-1.7\times 10^{-4}\mathrm{yr}^{-1}$ and $\beta\approx 0.5005$ with $M=0.21\mathrm{\leavevmode\nobreak\ M_{\odot}}$. The change of $\beta$ induced by mass transfer is $\langle\dot{\beta}/\beta\rangle_{\dot{M}}\approx 1.1\times 10^{-4}\mathrm{\leavevmode\nobreak\ yr}^{-1}$ with $\zeta=-1/3$. The bolometric luminosity of TDE is about $10^{44}\mathrm{\leavevmode\nobreak\ erg/s}$ corresponding to $\sim 1.6\dot{M}_{Edd}$. Assuming $e=0.95$, it predicts $\langle\dot{\beta}/\beta\rangle_{disk}\approx-3.5\times 10^{-4}\mathrm{\leavevmode\nobreak\ yr}^{-1}$ as shown in Fig. 6. The effects of disk drag and mass transfer on $\beta$ are in the same order of magnitude, suggesting that disk drag has the ability to stabilize mass transfer.
@@ 3.2.2 History and future of evolution
As the accretion rate decreases, drag force by the disk begins to dominate $\dot{\beta}$, and at that point the intensity of the QPEs begins to decrease. When the TDE rebrighten, the increased accretion rate led to a weaker drag force and thus an increase in the mass transfer. Then, when the accretion rate begins to decrease again, the effect of the drag force enhances and the intensity of QPEs will keep decreasing. After switching from radiation pressure dominance to gas pressure dominance in the inner region of the disk, the drag force decreases with decreasing accretion rate. When $\dot{\beta}$ is dominated by gravitational wave radiation, the mass transfer begins to enhance. When $\dot{\beta}$ is dominated by mass transfer, the unstable mass transfer leads to WD escape from the SMBH.
@@
3.2.3 Secular evolution of GSN 069
Due to the evolution on the accretion rate of the fallback disk in GSN 069, the secular evolution on mass transfer may experience complex changes. Fig. 9 show an example with possible parameters of GSN 069. Following the formula in Miniutti et al. (2023b), the accretion rate of each peak for TDE is described by
$\dot{M}=\dot{M}_{\text{peak }}\times\cases{e}^{-\left(t-t_{\text{peak }}\right)^{2}/2\sigma^{2}}&t\leq t_{\text{peak }}\\ {\left[\left(t-t_{\text{peak }}+t_{0}\right)/t_{0}\right]^{n}}t>t_{\text{peak }}{},$ (24)
where the parameters can be obtained by fitting the light curve (Miniutti et al., 2023b). For the first peak, $t_{0}=7.6$ years and $t_{\text{peak }}=600$ days, only the decay phase is calculated. For the second peak, $t_{0}=4.8$ year, $t_{\text{peak }}=3850$ days and $\sigma=400$ days. And $n$ is -9/4 for each peak, $\dot{M}_{\text{peak }}$ is 9 $\dot{M}_{Edd}$ for the first peak and 6.3 $\dot{M}_{Edd}$ for the second peak. The orbital parameters are same as them in Section 3.2.1, while the time is setting to $7$ years. Then the past and future evolution can be obtained by the same method in Section 2.3.
Figure 9: Possible evolution of mass transfer with the evolution of the accretion rate of the accretion disk. Solid line and dashed line represent the evolution of the mass transfer at the periastron and the accretion rate of the disk in GSN 069, respectively. See text for details of calculations.
It should be stressed that the results in Fig. 9 is very sensitive to the magnitude of parameters, mainly $\beta$ and $\dot{M}_{\text{peak }}$. The former affects the magnitude of $\dot{M}$ as it enlarges by a factor of $1/(\beta-\beta_{0})$ compared to the $\dot{\beta}$, as shown in Eq. 3. The latter, on the other hand, affects the magnitude of the disk drag, thus changing the moment of transition when $\beta$ changes from increasing to decreasing.
We do not show the evolution of the light curve directly, because the radiative processes are not considered here. If one assumes a positive correlation between the bolometric luminosity and the accretion mass, the disappearance of the QPEs in the early stages of the TDE and the lack of increase of its luminosity as well as its fading can be explained. That is, the early accreted mass increases rapidly and then starts to decrease as the disk drag increases.
However, the subsequent evolution is somewhat difficult to explain. Although the decrease in the accreted mass when the TDE flares again seems to explain the disappearance of the QPEs. But as pointed out by Miniutti et al. (2023a), the spectrum of QPEs need to be colder than it of the TDE to be undetectable, not just less luminous than the TDE. In addition, the second flare of the two QPEs that reappeared brightened, as expected in Fig. 9, yet the first flare did not brighten and could not be explained by our model.
These variations may be related to radiation processes of the QPEs. Lu & Quataert (2023) suggested that the QPEs is produced by the circularized shock of the accreted matter moving inside the disk. The peak temperature decreases during the second outburst of the TDE due to the decrease in disk density. This may explain the disappearance of QPEs. Thus the evolution of QPEs is also directly related to the evolution of accretion disk. The secular evolution on the light curve and spectrum of QPEs requires a further analysis combined with radiative processes in the future.
For the accretion model of the QPEs, the long-short behavior on the intensity and recurrence time is difficult to explain. King (2022) proposed that the short accretion timescale may cause the WD to fail to maintain hydrostatic equilibrium and oscillate in radius, which may explain this feature of QPEs. Such oscillations may be perturbed when the density of the surrounding gas changes, which may explain the disappearance of such features in QPEs before the second outburst of the TDE and the weaker QPE after this outburst that did not enhance as we expected. The evolution of such short timescales is beyond the scope of this paper, which focuses on long-term evolution.
@@ 4.3 On the other QPEs
The TDE in GSN 069 makes it unique. On the one hand, it may be the only source whose accretion disk is coplanar with the orbit of the stripped star. On the other hand, the accretion rate of its disk is dominated by the TDE and therefore evolves rapidly. These two points lead to a complicated orbital evolution, making the QPEs in GSN 069 an excellent source to study the interaction between the disk and the object moving inside it.
@@ 5 Summary
The TDE in GSN 069 may be a disruption of the common envelope before the WD was captured, so the orbiting WD is likely moving inside the fallback disk. The hydrodynamic drag force by the disk can effectively reduce $\beta$ and stabilize the mass transfer, which can explain the observed long term evolution in the intensity of QPEs. After the TDE re-brightened, the mass transfer increases due to the weakening of the drag force caused by the reduced density of the inner zone of the accretion disk.As the accretion rate of the fallback disk gradually decreases, the drag force by the disk will decrease until it is negligible. Then the mass transfer will start to be unstable eventually the WD will escape the SMBH.
The spectrum evolution of QPEsis also directly related to the evolution of the accretion disk. Future work will require further quantitative analysis of the evolution of mass transfer in combination with specific radiative processes. And continued observations of GSN 096 will help us to understand the interaction of the disk with the star inside it and to test our model.
@@ Acknowledgements.
Thanks Mengye Wang for helpful discussion. This work was supported by the National Key Research and Development Program of China (No. 2020YFC2201400) and National SKA Program of China (2020SKA0120300).